If the average of 3, 8 and x is equal to the average of 7 and x, what is the value of x?
Answer by mananth(16949)

3+8+x /3 = 7+x /2
..
11+x /3 =7+x /2
2(11+x)=3(7+x)
22+2x=21+3x
3x-2x=22-21
x=1